OpenStudy (anonymous):

what are the slope and y-intercept ? -9x + 10y =-9

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

m=9/10,b=-9/10 where m is the slope and b is the y-intercept

OpenStudy (triciaal):

rewrite the equation in the form y = m x + b the slope will be m the coefficient of x and b is the intercept

OpenStudy (anonymous):

how you do that

OpenStudy (anonymous):

hello

OpenStudy (triciaal):

if I told you 2 + 3 = 5 then you know if I ask what is 2 = you would do what is necessary to leave 2 on one side you would -3 and since you do to one side you would do to the other to keep the balance so 2 = 5-3 use the same principle and find what y =

OpenStudy (triciaal):

you isolate what you want and do what is necessary to keep the balance next you want the coefficient of the variable to be 1 for this problem you have 10 y and you need 1 y so divide by 10

OpenStudy (anonymous):

I got 9

OpenStudy (triciaal):

-9x + 10y =-9 add 9x to each side 10 y = -9 + 9x then divide by 10 y = -9/10 + 9/10* x standard to write the term in x before the constant so y = 9/10*x -9/10 slope = coefficient of x = 9/10 and intercept = constant = -9/10 are you ok with this now?

OpenStudy (triciaal):

@akon are you going to respond or should I skip your questions next time?

OpenStudy (anonymous):

thanks